OpenStudy (anonymous):

The width of a rectangle is fixed at 27 cm. What lengths will make the perimeter greater than 90 cm?

OpenStudy (pokemon23):

saifoo come to my question ;)

OpenStudy (saifoo.khan):

90-2(27) = 36. Lengths of both sides = 36. Length of one side = 36/2 = 18 So the length must be greater than 18cm.